OpenStudy (anonymous):

x/2x+2 - -2x/4x+4+2x-3/x+1

OpenStudy (anonymous):

\[\frac{ x }{ 2x+2 }-\frac{ -2x }{ 4x+4 }+\frac{ 2x-3 }{ x+1 }\]

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

least common multiple of the denominators, aka "common denominator" is \[4x+4\]

OpenStudy (misty1212):

\[\frac{ x }{ 2x+2 }-\frac{ -2x }{ 4x+4 }+\frac{ 2x-3 }{ x+1 }\] \[=\frac{ 2x }{ 4x+4 }-\frac{ -2x }{ 4x+4 }+\frac{4( 2x-3) }{ 4x+4 }\]

OpenStudy (anonymous):

my answer sheet gives the answer of three can you show me the steps to get to the answer

OpenStudy (triciaal):

you cannot solve without an equation check the original if you missed the =

OpenStudy (triciaal):

what @misty1212 did above is correct to simplify

OpenStudy (anonymous):

yes triciaal it didn't print right its = between the first two equations

OpenStudy (anonymous):

The following is a list of two expressions evaluated at x=7. The second expression is the first one, simplified. \[\left\{\frac{2 x-3}{x+1}+\frac{x}{2 x+2}+\frac{2 x}{4 x+4},3-\frac{6}{x+1}\right\}\text{/.}\, x\to 7 \]\[\left\{\frac{9}{4},\frac{9}{4}\right\} \]

OpenStudy (anonymous):

what misty did was right my print did print the paper out right

OpenStudy (triciaal):

I agree what Misty did was correct to simplify to solve the equation I have x = 3(x-2)